OpenStudy (anonymous):

Some help with these inequalities? Promise to give a medal! :) 1. y - 5 <= -8 2. y - 6 <= 2 3. y + 7 <= -5 4. y/5 > 5 5. x/5 >= -2

ganeshie8 (ganeshie8):

one step equations !

ganeshie8 (ganeshie8):

really oly one step we can do for each of them : isolate the variable by doing opposite operation

ganeshie8 (ganeshie8):

1. y - 5 <= -8 since y has -5 attached to it, add 5 both sides so that y gets isolated : y - 5 <= -8 + 5 + 5 y <= -3 DONE.

ganeshie8 (ganeshie8):

try next one, can u ? :)

OpenStudy (anonymous):

y - 6 <= 2 so i add 6 to both sides to get 8?

OpenStudy (anonymous):

and if there's a negative number do i reverse the inequality symbol or?

ganeshie8 (ganeshie8):

Yes !!!!

ganeshie8 (ganeshie8):

No, everything else remains same. just add 6 both sides.

OpenStudy (anonymous):

okay thank you, the rest of this quiz should be easy now. :) i definitely appreciate it!

ganeshie8 (ganeshie8):

np :) good luck !

ganeshie8 (ganeshie8):

5. x/5 >= -2 il do this ok.

ganeshie8 (ganeshie8):

5. x/5 >= -2 multiply 5 both sides x/5 >= -2 *5 *5 x >= -10

ganeshie8 (ganeshie8):

thats it, negative sign wont change anything here.

OpenStudy (anonymous):

awesome! thank you